# Rice-like complexity lower bounds for Boolean and uniform automata networks

### v1

@@ 6 Complexity lower bounds for $q$-uniform networks
Theorem 1.b, 1.d.
Let $q\geq 2$. For any $q$-non-trivial (resp. $q$-arborescent) MSO formula $\psi$, problem $\psi$-$q$-AN-dynamics (resp. $\psi$-$q$-NAN-dynamics) is ${\mathsf{NP}}$- or ${\mathsf{coNP}}$-hard.
@@
Again, the major difficulty is in the adaptation of Proposition 5 to $q$-uniform networks, i.e. to ensure that the pumping gives graphs whose sizes are powers of $q$. To this purpose we will employ Theorem 10 and the equivalent formulation of its condition in Theorem 19. This latter is expressed in terms of the coprime power of $q$ for $a$, but $a$ (the size of the pumped graph) is not known at this stage. In order to overcome this difficulty, we first fix adequate representatives for the equivalence classes of graphs, hence with a finite number of sizes. Afterwards, we will be able to get a value upper bounding the coprime power requirement of Theorem 19, to unlock the arithmetics with Theorem 10.
6.1 Equivalence classes representatives
Let $\psi$ (an MSO formula) and $k$ (an integer, equal to $2$ in the deterministic case) be fixed. Given a graph $G$ and a $k$-tree-decomposition $T$, we recall that $\mathcal{C}(v)$ (or $\mathcal{C}^{\psi}(v)$ in the deterministic case) is the equivalence class of $\mathcal{N}(v)$ for the relation $\sim_{k+1,\psi}$, where $\mathcal{N}(v)$ is a biboundaried graph with primary ports $B(v)$. Remark that since $T$ is a $k$-tree-decomposition, for any $v\in V(T)$, $\mathcal{C}(v)$ (or $\mathcal{C}^{\psi}(v)$) belongs to $\Sigma_{k+1,\psi}$. Most importantly, recall that $\Sigma_{k+1,\psi}$ is finite. In the deterministic case, $\Sigma_{3,\psi,\chi}$ is the analogous finite set of equivalence classes for both $\psi$ and $\chi$. Depending on whether we consider the deterministic or non-deterministic case, we will consider $\Sigma_{k+1,\psi,\chi}$ or $\Sigma_{k+1,\psi}$, so to combine the exposure, let us simply denote $\Sigma_{k+1}$ this set.
Let $\Sigma^{\prime}\subseteq\Sigma_{k+1}$ be such that for every class in $\Sigma^{\prime}$ there exists a representative $\sigma$, which is therefore a $(k+1)$-graph, such that:
  • $\sigma$ admits a $k$-tree-decomposition $T$, and
  • there is $v\in V(T)$ such that $\mathcal{N}(v)\subsetneq\sigma$ and $\mathcal{N}(v)\sim\sigma$.
@@ 6.1 Equivalence classes representatives
Let $s^{\prime}=|\Sigma^{\prime}|$, and denote $\{\sigma_{1},\dots,\sigma_{s^{\prime}}\}$ those representatives. The first condition will be fulfilled by our hypothesis on the non-triviality or arborescence of the formula $\psi$, and the second condition will correspond to the base case for pumping models (by gluing copies of $\mathcal{N}(v)$ to $\sigma$).
For all $i\in[s^{\prime}]$, let $\tau_{i}$ be a $k$-tree-decomposition of $\sigma_{i}$ whose root $r_{i}$ has a bag which corresponds to the primary ports of $\sigma_{i}$, hence $B(r_{i})=P_{1}(\sigma_{i})$. By construction, $\tau_{i}$ contains a node $t$ which is $\mathcal{N}$-different from the root and verifies $\mathcal{N}(t)\sim\sigma_{i}$. Remark that $k+1=|B(r_{i})|=|P_{1}(\sigma_{i})|$, because by convention all bags have size $k+1$. For all $i\in[s^{\prime}]$, denote:
$\tau^{\prime}_{i}=(\tau_{i}\setminus\mathcal{S}(t))\cup\{t\}\quad\text{and}\quad\sigma^{\prime}_{i}=\mathcal{N}_{\tau^{\prime}_{i}}(r_{i})$
the $(k+1)$-graph whose primary ports are the ports of $\sigma_{i}$ and secondary ports are the vertices in the bag of $t$. It holds that:
$\sigma_{i}=\sigma^{\prime}_{i}\oplus\mathcal{N}_{\tau_{i}}(t)\quad\text{and}\quad\tau_{i}=\tau^{\prime}_{i}\oplus\mathcal{S}_{\tau_{i}}(t).$
@@
Given that these representatives are fixed and a finite set, we will be able to upper bound the coprime power from Section 5 for the arithmetics of pumping $q$-uniform dynamics, independently of the initial model. Indeed, in the next construction we will substitute the pumped part (called $\tilde{G}_{1}$ so far) with its representative among $\{\sigma^{\prime}_{i}\}_{i\in[s^{\prime}]}$, because it will belong to $\Sigma^{\prime}$ by hypothesis.
@@ 6.2 Unbounded pumping
Let $x$ be an arbitrary integer, $\psi$ an MSO formula and an integer $k$. If $\psi$ has infinitely many $q$-uniform models of treewidth $k$, then there exist $\tilde{\Gamma}=\{\tilde{G}_{1},\tilde{G}_{2},\tilde{G}_{3}\}$ three $(k+1)$-graphs such that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)\models\psi$ for all $\ell\in\mathbb{N}$, and $P_{1}(\tilde{G}_{1})\cap P_{2}(\tilde{G}_{1})\neq V(\tilde{G}_{1})$. Moreover, $|\tilde{G}_{2}\oplus\tilde{G}_{3}|=q^{N}$ for some $N$ verifying $x\leq N$, and $\tilde{G}_{1}=\sigma^{\prime}_{i}$ for $i\in[s^{\prime}]$.
We will apply Proposition 20 for some $x$ related to the coprime powers of representatives $\{\sigma^{\prime}_{i}\}_{i\in[s^{\prime}]}$, but the following proof does not rely on any distinctive feature of coprime powers, therefore we state it for an arbitrary value $x$.
@@
As in the proof of Proposition 6, since $\psi$ has infinitely many $q$-uniform models admitting $k$-tree-decompositions, we consider a family $(M_{i})_{i\in\mathbb{N}}$ of such models and $(D_{i})_{i\in\mathbb{N}}$ a sequence of their $k$-tree-decompositions, such that each $D_{i}$ has at least one path of $\mathcal{N}$-length $i$. It holds for all $i\in\mathbb{N}$ that $|M_{i}|=q^{N}$ for some $N$, and we furthermore take them so that $x\leq N$. Without loss of generality we assume that every $M_{i}$ has at least $k+1$ vertices, and that all the bags of $D_{i}$ are of size $k+1$.
The rest of the proof is analogous to Proposition 6: $D_{i}$ are seen as $\Sigma_{k+1}$-labeled trees, and a path whose $\mathcal{N}$-length exceeds the size of $\Sigma_{k+1}$ has two nodes of the same equivalence class, within $\Sigma^{\prime}$. Denote $s=|\Sigma_{k+1}|$. Let us consider $D_{s+1}$ and $v,v^{\prime}$ those two nodes. With $i\in[s^{\prime}]$ such that $\sigma_{i}\sim\mathcal{N}(v)$, define:
$\tilde{T}_{2}=(D_{s+1}\setminus\mathcal{S}(v))\cup\{v\},\qquad\tilde{T}_{3}=\mathcal{S}(v),\qquad\tilde{T}_{1}=\tau^{\prime}_{i}.$
Figure 6: Construction of $\tilde{T}_{1},\tilde{T}_{2},\tilde{T}_{3}$ from $D_{s+1}$, in the proof of Proposition 20. The dotted part (to be pumped as $\tilde{T}_{1}$ between $\tilde{T}_{2}$ and $\tilde{T}_{3}$) is replaced with the tree-decomposition $\tau^{\prime}_{i}$ of $\sigma^{\prime}_{i}$ corresponding to representative $\sigma_{i}$ equivalent to $\mathcal{N}(v)$, with $i\in[s^{\prime}]$.
See Figure 6. With appropriate ports taken as the bags of $v$ and $v^{\prime}$, we have $\tilde{T}_{2}\oplus\tilde{T}_{3}=D_{s+1}$. By [10, Lemma 14] we deduce that $\Lambda^{\mathcal{T}}(2\cdot 1^{\ell}\cdot 3)$ is positive for $\psi$ for all $\ell\in\mathbb{N}$, and by Remark 2 there exist $\tilde{\Gamma}=\{\tilde{G}_{1},\tilde{G}_{2},\tilde{G}_{3}\}$ associated respectively to $\tilde{T}_{1},\tilde{T}_{2},\tilde{T}_{3}$ such that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ is a model of $\psi$ for all $\ell\in\mathbb{N}$. By construction we still have $P_{1}(\tilde{G}_{1})\cap P_{2}(\tilde{G}_{1})\neq V(\tilde{G}_{1})$, and furthermore $\tilde{G}_{2}\oplus\tilde{G}_{3}=M_{s+1}$ hence it is a $q$-uniform model of size greater than $q^{x}$, and finaly $\tilde{G}_{1}=\sigma^{\prime}_{i}$ (in the construction on tree-decompositions, the pumped part is replaced by $\tau^{\prime}_{i}$ corresponding to the graph $\sigma^{\prime}_{i}$). ∎
@@ 6.3 Assembling $q$-uniform networks
We prove Theorem 1.b, 1.d for any $q\geq 2$, based on arithmetical considerations generalizing the case $q=2$ presented in Subsection 5.2. The proof structure consists in applying Proposition 20 to obtain three $(k+1)$-graphs, then incorporte the saturating graph $\Omega_{m}$, and finaly reduce from SAT or UNSAT by constructing automata networks whose sizes are powers of $q$, thanks to the arithmetics from Theorem 10.
@@
One may remark from the condition of Proposition 20 on the size of $\tilde{G}_{2}\oplus\tilde{G}_{3}$ that the case $\ell=0$ will be important, and justifies its consideration in Section 5. It is required by our reasoning, in order to control the size of the initial solution to ${\mathcal{P}_{q}(a,b)}$.
@@
The other key ingredient consists in replacing the pumped part by its representative among $\{\sigma^{\prime}_{i}\}_{i\in[s^{\prime}]}$, which allows to bound the value of $a$ and apply Proposition 20 with a value of $x$ suitable for the subsequent use of the equivalent condition given by Theorem 19 for the application of Theorem 10.
Proof of Theorem 1.b, 1.d.
@@
The proof structure is analogous to Theorem 1 for $q=2$, with additionnal considerations to apply the arithmetics from Section 5. The deterministic case is handled with the addition of $\chi$ as in Section 4, in order to have graphs of out-degree $1$ all the way.
@@
Let $q\geq 2$, and $\psi$ be a $q$-arborescent (or $q$-non-trivial) MSO formula of rank $m$. Without loss of generality, assume that the saturating graph $\Omega_{m}$ from Proposition 2 (or Corollary 3) is a counter-model of $\psi$, and let us build a reduction from UNSAT (otherwise consider $\neg\psi$ instead of $\psi$, and a reduction from SAT instead of UNSAT).
@@
Before calling Proposition 20, we setup an apropriate value of $x$ to ensure that the condition of Theorem 19 will be verified. It is related to the value of $a$ which is the size of the glued part, i.e. $\sigma^{\prime}_{i}$ from the Proposition. Recall that $\oplus$ merges the ports, of size $k$ in our context. Let $\eta$ be a value greater or equal to the maximum coprime power of $q$ for $\{|\sigma^{\prime}_{i}|-k\mid i\in[s^{\prime}]\}$, and also greater or equal to $\max\{\log_{q}(|\sigma^{\prime}_{i}|-k)\mid i\in[s^{\prime}]\}$.
@@
Applying Proposition 20 for $x=\eta$, we obtain three $k$-graphs $\tilde{\Gamma}=\{\tilde{G}_{1},\tilde{G_{2}},\tilde{G}_{3}\}$, with $\tilde{G}_{1}=\sigma^{\prime}_{i}$ for some $i\in[s^{\prime}]$, such that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)\models\psi$ for all $\ell\in\mathbb{N}$. From $P_{1}(\tilde{G}_{1})\cap P_{2}(\tilde{G}_{1})\neq V(\tilde{G}_{1})$ we deduce the existence of $(a,b)\in\mathbb{N}_{+}\times\mathbb{N}$ with:
  • $b=|\tilde{G}_{2}\oplus\tilde{G}_{3}|=q^{N}$ for some $N\geq\eta$,
  • $a=|\tilde{G}_{1}|-k=|\sigma^{\prime}_{i}|-k$.
@@
We have $|\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)|=a\ell+b$, which is a power of $q$ for $\ell=0$, i.e. we have the initial solution $(0,N)$ to ${\mathcal{P}_{q}(a,b)}$. Since this solution verifies that the coprime power of $q$ for $a$ is smaller than $N$, by Theorem 19 there exists a couple $(\mu,\kappa)\in\mathbb{N}_{+}\times\mathbb{N}_{+}$ such that we can apply Theorem 10, and deduce that ${\mathcal{P}_{q}(a,b)}$ contains a geometric sequence of integers. More precisely, from the proof of Lemma 15, in $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ we have graphs of size $q^{N+\ell^{\prime}\mu}=bq^{\ell^{\prime}\mu}$ for all $\ell^{\prime}\in\mathbb{N}$, which correspond to $\ell=\frac{bq^{\ell^{\prime}\mu}-b}{a}$ being an integer (the number of copies of $\tilde{G}_{1}$).
@@
Before continuing, let us consider the fourth graph. Let $\tilde{G_{0}}=\Omega_{m}\sqcup\tilde{G_{1}}$. Let $\alpha$ be the least power of $q$ such that $G_{1}=\bigoplus_{i=1}^{\alpha}\tilde{G}_{1}$ verifies $|G_{1}|\geq|\tilde{G}_{0}|$, and let $G_{0}$ be the disjoint union of $\tilde{G}_{0}$ and $|G_{1}|-|\tilde{G}_{0}|$ fixed points, so that $|G_{0}|=|G_{1}|$. Finaly, let $G_{2}=\tilde{G}_{2}$ and $G_{3}=\tilde{G}_{3}$. The $k$-graphs $\Gamma=\{G_{0},G_{1},G_{2},G_{3}\}$ verify Conditions 1–3 of Proposition 4, as usual. Furthermore, $G_{1}$ is made of $\alpha$ copies of $\tilde{G}_{1}$.
@@
Given a propositional formula $S$ on $s$ variables, we need at least $2^{s}$ copies of $G_{0}$ or $G_{1}$ in order to implement the reduction from UNSAT, where for each valuation the circuit of the AN (resp. NAN) builds a copy of $G_{0}$ if $S$ is satisfied, and a copy of $G_{1}$ otherwise. Recall that $G_{1}$ is made of $\alpha$ copies of $\tilde{G}_{1}$ and $|G_{0}|=|G_{1}|$, that is, for $|w|=2^{s}$ we have $|\Delta^{\Gamma}(2\cdot w\cdot 3)|=a\alpha 2^{s}+b$. The goal is now to find a suitable padding by $L(s)$ copies of $\tilde{G}_{1}$, i.e. such that:
@@
is a power of $q$, with $|w|=2^{s}$, $G_{4}=\tilde{G}_{1}$ and $\Gamma^{\prime}=\{G_{0},G_{1},G_{2},G_{3},G_{4}\}$. Then the reduction is as usual: the graph $\Delta^{\Gamma^{\prime}}(2\cdot w\cdot 4^{L(s)}\cdot 3)$ is a model of $\psi$ if and only if $w$ does not contain letter 00 (in this case it has only copies of $\tilde{G}_{1}$). Remember that the only non-constant part is $S$ (and $s$).
@@
Therefore, we can build a Boolean NAN (or AN) with
@@
$\alpha 2^{s}+L(s)=\frac{bq^{\mu(s+1)}\alpha^{\mu}-b}{a}$ copies of $\tilde{G}_{1}$
(some copies of $\tilde{G}_{1}$ form copies of $G_{1}$, or are replaced by a copy of $G_{0}$ when $S$ is satisfied). Recall that $a$, $b$, $\mu$ and $\alpha$ are constants depending only on $\psi$. ∎

### v2

@@ 6 Complexity lower bounds for $q$-uniform networks
Theorem 1.b,d.
Let $q\geq 2$. For any $q$-non-trivial (resp. $q$-cw-non-trivial) MSO formula $\psi$, problem $\psi$-$q$-AN-dynamics (resp. $\psi$-$q$-NAN-dynamics) is $\mathsf{NP}$}- or $\mathsf{coNP}$}-hard.
@@
Again, the major difficulty is in the adaptation of Proposition 5 to $q$-uniform networks, i.e. to ensure that the pumping gives graphs whose sizes are powers of $q$. To this purpose, we will employ Theorem 13 and the equivalent formulation of its condition in Theorem 22. This latter is expressed in terms of the coprime power of $q$ for $a$, but $a$ (the size of the pumped graph) is not known at this stage. In order to overcome this issue, we first fix adequate representatives for each MSO type of graph, hence with a finite number of sizes. Afterwards, we will be able to get a value upper bounding the coprime power requirement of Theorem 22, to unlock the arithmetics with Theorem 13.
6.1 Type representatives
Recall that $T_{m}$ is the set of realized types of quantifier rank $m$ and it is finite. In the following, we denote $t(\mathcal{C})$ the type of the graph generated by $\mathcal{C}$, the latter being written $G_{\mathcal{C}}$. Moreover, for a node $v$ in $\mathcal{C}$ we still denote $\mathcal{C}(v)$ the clique-decomposition composed by $v$ and all its descendant (the subtree whose root is $v$), and $t(v)=t(\mathcal{C}(v))$.
Let $T^{\prime}_{m}\subseteq T_{m}$ be the set of realized types of quantifier rank $m$ such that, for each type $t\in T^{\prime}_{m}$, there exists a representative clique-decomposition $\tau$ such that:
  • $t(\tau)=t$,
  • there is a node $v\in\tau$ which is not the root such that $t(\tau(v))=t$.
@@ 6.1 Type representatives
Let $s^{\prime}=|T^{\prime}_{m}|$, and denote $\{\tau_{1},\dots,\tau_{s^{\prime}}\}$ those representatives. The first condition helps identifying representatives, and the second condition will correspond to the base case for pumping models (by gluing copies of $\tau(v)$ inside $\tau$), which will be given by our hypothesis on the non-triviality or the cw-non-triviality of the formula $\psi$. For all $i\in[s^{\prime}]$, let $\tau^{\prime}_{i}$ be $\tau_{i}$ where $\tau_{i}(u)$ has been removed and $u$ has been replaced by $\square$. Hence $\tau_{i}=\tau^{\prime}_{i}\oplus\tau(u)$.
@@
Given that these representatives are fixed and form a finite set, we will be able to upper bound the coprime power from Section 5 for the arithmetics of pumping $q$-uniform dynamics, independently of the initial model. Indeed, in the next construction we will substitute the pumped part (called $\tilde{\mathcal{C}}_{1}$ so far) with its representative among $\{\tau^{\prime}_{i}\}_{i\in[s^{\prime}]}$, because its type will belong to $T^{\prime}_{m}$ by hypothesis.
@@ 6.2 Unbounded pumping
Let $x$ be an arbitrary integer, $\psi$ an MSO formula and an integer $k$. If $\psi$ has infinitely many $q$-uniform models of cliquewidth at most $k$, then there exist $\tilde{\Gamma}=\{\tilde{\mathcal{C}}_{1},\tilde{\mathcal{C}}_{2},\tilde{\mathcal{C}}_{3}\}$ three clique-decomposition of width at most $k$ such that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)\models\psi$ for all $\ell\in\mathbb{N}$. Moreover, $|\tilde{\mathcal{C}}_{2}\oplus\tilde{\mathcal{C}}_{3}|=q^{N}$ for $N$ verifying $x\leq N$, and $\tilde{\mathcal{C}}_{1}=\tau^{\prime}_{i}$ for $i\in[s^{\prime}]$.
We will apply Proposition 23 for some $x$ related to the coprime powers of representatives $\{\tau^{\prime}_{i}\}_{i\in[s^{\prime}]}$, but the following proof does not rely on any distinctive feature of coprime powers, therefore we state it for an arbitrary value $x$.
@@
As in the proof of Proposition 9, since $\psi$ has infinitely many $q$-uniform models, and since there are finitely many types, there must be a type $t_{0}\in T_{m}$ which contains $\psi$ and is realized by infinitely many $q$-uniform graphs of cliquewidth at most $k$. Let $\mathcal{C}$ be a clique-decomposition of cliquewidth at most $k$ of height at least $|T_{m}|+1$, such that $t(\mathcal{C})=t_{0}$ and $|\mathcal{C}=q^{N}$ for some $N\in\mathbb{N}$. Without loss of generality we assume that $\mathcal{C}$ has at least $q^{x}$ constant nodes (otherwise, we would just have to consider a bigger $q$-uniform model).
Any path of length at least $|T_{m}|+1$ in a clique-decomposition contains at least two different nodes $v$ and $v^{\prime}$ such that $t(\mathcal{C}(v))=t(\mathcal{C}(v^{\prime}))$. Let $v$ and $v^{\prime}$ be two such nodes in a path of length at least $|T_{m}|+1$ in $\mathcal{C}$. We assume without loss of generality that $v^{\prime}$ is a descendant of $v$. Because of the characteristics of $\mathcal{C}(v)$, there is a representative $\tau_{i}\in T_{m}^{\prime}$ such that $t(\mathcal{C}(v))=t(\tau_{i})$, we denote it $t$.
Let $\tilde{\mathcal{C}_{3}}=\mathcal{C}(v)$, $\tilde{\mathcal{C}_{1}}=\tau_{i}^{\prime}$, and $\tilde{\mathcal{C}_{2}}$ be $\mathcal{C}$ where $\mathcal{C}(v)$ has been removed and $v$ has been replaced by $\square$. These definitions are such that $\mathcal{C}=\tilde{\mathcal{C}_{2}}\oplus\tilde{\mathcal{C}_{3}}$, see Figure 7. Remark that $\mathcal{C}_{1}$ is not empty since $v\neq v^{\prime}$ (Remark 2).
Figure 7: Construction of $\tilde{C}_{1},\tilde{C}_{2},\tilde{C}_{3}$ in the proof of Proposition 23. The dotted part (to be pumped as $\tilde{C}_{1}$ between $\tilde{C}_{2}$ and $\tilde{C}_{3}$) is replaced with the clique-decomposition $\tau^{\prime}_{i}$ corresponding to representative $\tau_{i}$ equivalent to $\mathcal{C}(v)$, with $i\in[s^{\prime}]$.
Recall that $t(\tilde{\mathcal{C}_{3}})=t(\mathcal{C}(v))=t(\mathcal{C}(v^{\prime}))$, and since $t(\mathcal{C}(v))=t(\tau_{i})$ all these four clique-decompositions have the same type $t$. Moreover, let $u$ the node in $\tau_{i}$ such that $\tau_{i}=\tau^{\prime}_{i}\oplus\tau_{i}(u)$ and $t(\tau_{i})=t(\tau_{i}(u))$. Hence, considering the map $\Lambda_{\tau^{\prime}_{i}}$ from Lemma 6, we have $\Lambda_{\tau^{\prime}_{i}}(t)=t$. Thus, gluing any of the four clique-decomposition having type $t$ to $\tau^{\prime}_{i}$ will give a graph of type $t$. In particular, since $\tilde{\mathcal{C}_{1}}=\tau^{\prime}_{i}$, the type of $\tilde{\mathcal{C}_{1}}\oplus\dots\oplus\tilde{\mathcal{C}_{1}}\oplus\tilde{\mathcal{C}_{3}}$ is $t$, whatever the number of repetitions of $\tilde{\mathcal{C}_{1}}$. Finally, since $\mathcal{C}=\tilde{\mathcal{C}_{2}}\oplus\tilde{\mathcal{C}_{3}}$ has type $t_{0}$, we have $\Lambda_{\tilde{\mathcal{C}_{2}}}(t)=t_{0}$ and the type of $\tilde{\mathcal{C}_{2}}\oplus\tilde{\mathcal{C}_{1}}\oplus\dots\oplus\tilde{\mathcal{C}_{1}}\oplus\tilde{\mathcal{C}_{3}}$ is always $t_{0}$, which concludes the proof because $\phi\in t_{0}$. ∎
@@ 6.3 Assembling $q$-uniform networks
We prove Theorem 1.b,d for any $q\geq 2$, based on arithmetical considerations generalizing the case $q=2$ presented in Subsection 5.2. The proof structure consists in applying Proposition 23 to obtain three clique-decompositions of width at most $k$, then incorporate the saturating graph $\Omega_{m}$, and finally reduce from SAT or UNSAT by constructing automata networks whose sizes are powers of $q$, thanks to the arithmetics from Theorem 13.
@@
One may remark from the condition of Proposition 23 on the size of $\tilde{\mathcal{C}}_{2}\oplus\tilde{\mathcal{C}}_{3}$ that the case $\ell=0$ will be important, and justifies its consideration in Section 5. It is required by our reasoning, in order to control the size of the initial solution to $\mathcal{P}_{q}(a,b)$}.
@@
The other key ingredient consists in replacing the pumped part by its representative among $\{\tau^{\prime}_{i}\}_{i\in[s^{\prime}]}$, which allows to bound the value of $a$ and apply Proposition 23 with a value of $x$ suitable for the subsequent use of the equivalent condition given by Theorem 22 for the application of Theorem 13.
Proof of Theorem 1.b,d.
@@
The proof structure is analogous to Theorem 1 for $q=2$, with additional considerations to apply the arithmetics from Section 5. The deterministic case is handled with the addition of $\chi$ as in Section 4, in order to have graphs of out-degree $1$ all the way.
@@
Let $q\geq 2$, and $\psi$ be a $q$-cw-non-trivial (or $q$-non-trivial) MSO formula of rank $m$. Without loss of generality, assume that the saturating graph $\Omega_{m}$ from Proposition 3 (or Corollary 4) is a counter-model of $\psi$, and let us build a reduction from UNSAT (otherwise consider $\neg\psi$ instead of $\psi$, and a reduction from SAT instead of UNSAT).
@@
Before calling Proposition 23, we setup an appropriate value of $x$ to ensure that the condition of Theorem 22 will be verified. It is related to the value of $a$ which is the size of the glued part, i.e. $\tau^{\prime}_{i}$ from the Proposition. Let $\eta$ be a value greater or equal to the maximum coprime power of $q$ for $\{|\tau^{\prime}_{i}|\mid i\in[s^{\prime}]\}$, and also greater or equal to $\max\{\log_{q}(|\tau^{\prime}_{i}|)\mid i\in[s^{\prime}]\}$.
@@
Applying Proposition 23 for $x=\eta$, we obtain three $k$-clique-decompositions $\tilde{\Gamma}=\{\tilde{\mathcal{C}}_{1},\tilde{\mathcal{C}_{2}},\tilde{\mathcal{C}}_{3}\}$, with $\tilde{\mathcal{C}}_{1}=\tau^{\prime}_{i}$ for some $i\in[s^{\prime}]$, such that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)\models\psi$ for all $\ell\in\mathbb{N}$. We deduce the existence of $(a,b)\in\mathbb{N}_{+}\times\mathbb{N}$ with:
  • $b=|\tilde{\mathcal{C}}_{2}\oplus\tilde{\mathcal{C}}_{3}|=q^{N}$ for some $N\geq\eta$,
  • $a=|\tilde{\mathcal{C}}_{1}|=|\tau^{\prime}_{i}|$.
@@
We have $|\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)|=a\ell+b$, which is a power of $q$ for $\ell=0$, i.e. we have the initial solution $(0,N)$ to $\mathcal{P}_{q}(a,b)$}. Since this solution verifies that the coprime power of $q$ for $a$ is smaller than $N$, by Theorem 22 there exists a couple $(\mu,\kappa)\in\mathbb{N}_{+}\times\mathbb{N}_{+}$ such that we can apply Theorem 13, and deduce that $\mathcal{P}_{q}(a,b)$} contains a geometric sequence of integers. More precisely, from the proof of Lemma 18, in $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ we have graphs of size $q^{N+\ell^{\prime}\mu}=bq^{\ell^{\prime}\mu}$ for all $\ell^{\prime}\in\mathbb{N}$, which correspond to $\ell=\frac{bq^{\ell^{\prime}\mu}-b}{a}$ being an integer (the number of copies of $\tilde{\mathcal{C}}_{1}$).
@@
Before continuing, let us consider the fourth graph. Let $\alpha$ be the least power of $q$ such that $\mathcal{C}_{1}=\bigoplus_{i=1}^{\alpha}\tilde{\mathcal{C}}_{1}$ verifies $|\mathcal{C}_{1}|\geq|\Omega_{m}|$, and let $\mathcal{C}_{0}$ be the disjoint union of $\mathcal{C}_{m}$ (a clique-decomposition of $\Omega_{m}$) and $|\mathcal{C}_{1}|-|\Omega_{m}|$ constant nodes, so that $|\mathcal{C}_{0}|=|\mathcal{C}_{1}|$. Finally, let $\mathcal{C}_{2}=\tilde{\mathcal{C}}_{2}$ and $\mathcal{C}_{3}=\tilde{\mathcal{C}}_{3}$. The clique-decompositions $\Gamma=\{\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ verify Conditions (i)–(iii) of Proposition 7. Furthermore, $\mathcal{C}_{1}$ is made of $\alpha$ copies of $\tilde{\mathcal{C}}_{1}$.
@@
Given a propositional formula $S$ on $s$ variables, we need at least $2^{s}$ copies of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$ in order to implement the reduction from UNSAT, where for each valuation the circuit of the AN (resp. NAN) builds a copy of the graph generated by $\mathcal{C}_{0}$ if $S$ is satisfied, and a copy of the graph generated by $\mathcal{C}_{1}$ otherwise. Recall that $\mathcal{C}_{1}$ is made of $\alpha$ copies of $\tilde{\mathcal{C}}_{1}$ and $|\mathcal{C}_{0}|=|\mathcal{C}_{1}|$, that is, for $|w|=2^{s}$ we have $|\Delta^{\Gamma}(2\cdot w\cdot 3)|=a\alpha 2^{s}+b$. The goal is now to find a suitable padding by $L(s)$ copies of $\tilde{\mathcal{C}}_{1}$, i.e. such that:
@@
is a power of $q$, with $|w|=2^{s}$, $\mathcal{C}_{4}=\tilde{\mathcal{C}}_{1}$ and $\Gamma^{\prime}=\{\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3},\mathcal{C}_{4}\}$. Then the reduction is as usual: the graph generated by $\Delta^{\Gamma^{\prime}}(2\cdot w\cdot 4^{L(s)}\cdot 3)$ is a model of $\psi$ if and only if $w$ does not contain letter 0 (in this case it has only copies of $\tilde{\mathcal{C}}_{1}$). Remember that the only non-constant part is $S$ (and $s$).
@@
Therefore, we can build the circuit of a $q$-uniform NAN (or AN) with
@@
$\alpha 2^{s}+L(s)=\frac{bq^{\mu(s+1)}\alpha^{\mu}-b}{a}$ copies of $\tilde{\mathcal{C}}_{1}$
(some copies of $\tilde{\mathcal{C}}_{1}$ form copies of $\mathcal{C}_{1}$, or are replaced by a copy of $\mathcal{C}_{0}$ when $S$ is satisfied). Recall that $a$, $b$, $\mu$ and $\alpha$ are constants depending only on $\psi$. ∎